OpenStudy (anonymous):

19. A fan club comprising girls and boys has a total of 160 members. If 32 more girls join the club, the percentage of girls will increase by 5%. Find the number of boys in the club. Please show all workings clearly.

OpenStudy (mertsj):

b+g=160 (g+32)/192=(g/160) +.05

OpenStudy (mertsj):

Solve the second equation and you will get g = 112 Therefore the number of boys = 160-112 or 48